OpenStudy (anonymous):

simplify the radical expression sqrt3+4sqrt3 simplify the radical expression 2sqrt6+3sqrt96 simplify the radical expression (3+sqrt6)(3-sqrt6) solve the equation 7=sqrtx +2

OpenStudy (anonymous):

please help!!

OpenStudy (kc_kennylau):

\[\sqrt3+4\sqrt3=5\sqrt3\] \[2\sqrt6+3\sqrt{96}=2\sqrt6+3*4\sqrt6=14\sqrt6\] \[(3+\sqrt6)(3-\sqrt6)=3^2-\sqrt6^2=9-6=3\] \[7=\sqrt{x}+2\] \[5=\sqrt{x}\] \[x=25\]

OpenStudy (anonymous):

Help with more?

OpenStudy (kc_kennylau):

more=?

OpenStudy (anonymous):

solve for x. identify any extraneous solutions. sqrtx=-3 what is the domain of the function? y=sqrt5x-10